### v1

# Iterations of mean-type mappings and their convergence

Abstract.
We define so-called residual means, which have a Taylor expansion of the form $M(x)=\bar{x}+\tfrac{1}{2}\xi_{M}(\bar{x})\operatorname{Var}(x)+o(\|x-\bar{x}\|^{\alpha})$ for some $\alpha>2$ and a single-variable function $\xi_{M}$ ($\bar{x}$ stands for the arithmetic mean of the vector $x$), and show that all symmetric means which are three times continuously differentiable are residual. We also calculate the value of residuum for quasideviation means and a few subclasses of this family.
Later, we apply it to establish the limit of the sequence $\big{(}\frac{\text{Var}\ {\bf M}^{n+1}(x)}{(\text{Var}\ {\bf M}^{n}(x))^{2}}\big{)}_{n=1}^{\infty}$, where ${\bf M}\colon I^{p}\to I^{p}$ is a mean-type mapping consisting of $p$-variable residual means on an interval $I$, and $x\in I^{p}$ is a nonconstant vector.
Key words and phrases:
Gaussian iteration process, invariant means, quasideviation mean, $p$-variable means, mean-type mapping, Arrow-Pratt index
2010 Mathematics Subject Classification:
26E60, 26B15, 26A18, 39B12, 41A58
1. Introduction
Various aspects of means have been continuously studied for more than a hundred years. There is a wide variety of problems related to them, such as comparability, equality between means, invariance, etc. The value of means close to the diagonal (that is, the set where all entries are equal to each other) plays an important role in this theory. Let us just recall a classical result stating that if two quasiarithmetic means (see section 4.4 for the definition) are comparable in the vicinity of the diagonal, then they are comparable on the whole domain (see, for example, Hardy-Littlewood-Polyá [18]). Moreover, under the additional smoothness assumptions, for a given quasiarithmetic mean generated by a function $f$, we can define the related single variable function $\tfrac{f^{\prime\prime}}{f^{\prime}}$ (so-called Arrow-Pratt index) and show that the comparability of two quasiarithmetic means and their Arrow-Pratt indexes are equivalent (this fact is implied by Jensen’s inequality; see Mikusiński [31] for the most general wording of this statement).
Furthermore, it is known that iterations of mean-type mapping consisting of continuous strict means converge to some point on the diagonal (see, for example, [7, Theorem 8.8]). In 2018 it was proved [33] that in the case when all means are quasiarithmetic, the precise speed of convergence is related to the variance of Arrow-Pratt indexes of all these means. The precise result reads as follows.
Theorem A.
Let $p\in\mathbb{N}$, $I$ be an interval, $f_{1},\dots,f_{p}\colon I\to\mathbb{R}$ be strictly monotone, twice continuously differentiable functions such that each $f_{i}^{\prime\prime}$ is locally Lipschitz. Define the mean-type mapping $\mathbf{M}\colon I^{p}\to I^{p}$ by $\mathbf{M}(x):=(\mathscr{A}_{f_{1}}(x),\dots,\mathscr{A}_{f_{p}}(x))$, where $\mathscr{A}_{f_{i}}$ is the quasiarithmetic mean generated by $f_{i}$.
For all $x\in I^{p}$ either $\mathbf{M}^{n}(x)$ is constant for some $n\in\mathbb{N}$ or
$\lim_{n\to\infty}\frac{\operatorname{Var}\mathbf{M}^{n+1}(x)}{\big{(}\operatorname{Var}\mathbf{M}^{n}(x)\big{)}^{2}}=\frac{1}{4}\operatorname{Var}\bigg{(}\frac{f_{1}^{\prime\prime}(m)}{f_{1}^{\prime}(m)},\dots,\frac{f_{p}^{\prime\prime}(m)}{f_{p}^{\prime}(m)}\bigg{)},$
where $m$ is the value of the unique $\mathbf{M}$-invariant mean at $x$.
This theorem supports the idea of searching the generalization of the Arrow-Pratt index for other families of means. The comparability property will no longer be valid since the comparability of means is not localizable in general, there is however the hope to reprove Theorem A for the broader class of means. Our motivation comes from the theorem of Borwein-Borwein
Theorem B ([7], Theorem 8.8).
Let $p\in\mathbb{N}$, $I$ be an interval, $M_{1}\dots,M_{p}\colon I^{p}\to I$ be a strict means, $x\in I^{p}$, and $\mathbf{M}:=(M_{1},\dots,M_{p})\colon I^{p}\to I^{p}$. If all $M_{i}$-s are continuously differentiable and $\mathbf{M}^{n}(x)$ is nonconstant for all $n\in\mathbb{N}$ then
$\lim_{n\to\infty}\frac{\max\mathbf{M}^{n+1}(x)-\min\mathbf{M}^{n+1}(x)}{\max\mathbf{M}^{n}(x)-\min\mathbf{M}^{n}(x)}=0.$
If the means are twice continuously differentiable, convergence in the Gaussian iteration is quadratic (uniformly on compact subsets).
Based on Theorem B, the most natural assumption in a potential generalization of Theorem A would be the twice continuous differentiability of the means. Regretfully, our assumptions are more restrictive. We will deal with this in the next section. We define so-called >>residual<< means and show that all means which are three times continuously differentiable are residual (section 2). Then the Arrow-Pratt index corresponds to the >>residuum<< of a mean. Later, after some auxiliary results, we generalize Theorem A for mean-type mappings consisting of residual means (section 3). In the final section, we establish the value of residuum for quasideviation means and a few subclasses of this family.
2. Approximate values of means
This section, in a rough sketch, is devoted to estimating the value of $p$-variable $\mathcal{C}^{2}$ means $M\colon I^{p}\to I$ close to the diagonal (Lemma 2.1). Prior to this, let us formally introduce a few preliminary definitions.
Throughout this note $p\in\mathbb{N}$ and $I$ is an open subinterval of $\mathbb{R}$. Recall that a $p$-variable mean on $I$ is an arbitrary function $M\colon I^{p}\to I$ satisfying the mean property, that is
$\min(x_{1},\dots,x_{p})\leq M(x_{1},\dots,x_{p})\leq\max(x_{1},\dots,x_{p})\text{ for all }x_{1},\dots,x_{p}\in I.$
Quite often we additionally assume that $M$ is symmetric, which means that $M(x\circ\sigma)=M(x)$ for all $x\in I^{p}$ and a permutation $\sigma\colon\{1,\dots,p\}\to\{1,\dots,p\}$.
From now on let ${\bf 1}$ stands for the vector of the suitable dimension with all elements equal to one. Observe that the mean property forces the values of the mean on the diagonal
$\Delta_{p}(I):=\{(x_{1},\dots,x_{p})\in I^{p}\colon x_{1}=\cdots=x_{p}\}=\{x{\bf 1}\colon x\in I\}.$
More precisely, $M(x{\bf 1})=x$ for all $x\in I$ (this property alone is known as reflexivity).
We are heading towards the lemma, which establishes the value of a mean close to the diagonal. Later, we introduce the family of means which satisfies the stronger version of the inequality proved in this lemma (residual means). Then, in Theorem 2.3, we show that all means which are $\mathcal{C}^{3}$ close to the diagonal meet this property. The motivation for such an approach will be clarified in the next section.
For a vector $s\in\mathbb{R}^{p}$ let us introduce three classical, brief notations
${\bf E}s:=\frac{1}{p}\sum_{i=1}^{p}s_{i};\qquad{\bf E}s^{2}:=\frac{1}{p}\sum_{i=1}^{p}s_{i}^{2};\qquad\operatorname{Var}(s):={\bf E}s^{2}-({\bf E}s)^{2}.$
Now we are ready to proceed to the first lemma.
Lemma 2.1.
Let $M\colon I^{p}\to I$ be a symmetric mean of class $\mathcal{C}^{2}$ in the neighborhood of the diagonal $\Delta_{p}(I)$. There exists a continuous function $\xi_{M}\colon I\to I$ such that for all $x\in I$ and $s\in\mathbb{R}^{p}$ we have
(2.1) $M(x{\bf 1}+ts)$ $=x+t{\bf E}s+\frac{t^{2}}{2}\xi_{M}(x)\operatorname{Var}(s)+o(t^{2})\quad\text{ for }t\approx 0.$
Moreover, for all $x\in I$,
(2.2) $\xi_{M}(x)=-p^{2}\partial_{1}\partial_{2}M(x{\bf 1});\qquad\xi_{M}(x)=\frac{p^{2}}{p-1}\partial_{1}^{2}M(x{\bf 1}).$
Proof.
Fix $x\in I$ and $s\in\mathbb{R}^{p}$ arbitrarily. First, we use Taylor’s theorem. For $t$ close to zero, we get
(2.3) $M(x{\bf 1}+ts)$ $=M(x{\bf 1})+t\sum_{i=1}^{p}\partial_{i}M(x{\bf 1})s_{i}+\frac{t^{2}}{2}\sum_{i=1}^{p}\sum_{j=1}^{p}\partial_{i}\partial_{j}M(x{\bf 1})s_{i}s_{j}+o(t^{2})$
Since $M$ is symmetric, we know that the value of the partial derivatives $(\partial_{i}M)_{i=1}^{p}$ does not depend on $i$. The same holds for second partial derivatives (separately pure and mixed). Therefore, for the sake of brevity, let us define $\alpha,\beta,\gamma\colon I\to\mathbb{R}$ by
$\alpha(x):=\partial_{1}^{2}M(x{\bf 1});\quad\beta(x):=\partial_{1}\partial_{2}M(x{\bf 1});\quad\gamma(x):=\partial_{1}M(x{\bf 1}).$
Since $M$ is reflexive, we have $M(x{\bf 1})=x$ and $M((x+t){\bf 1})=x+t$, thus
$x+t=M\big{(}(x+t){\bf 1}\big{)}=M(x{\bf 1})+t\sum_{i=1}^{p}\partial_{i}M(x{\bf 1})=x+tp\gamma(x)\text{ for all }x\in I,$
which proves $\gamma\equiv\frac{1}{p}$. Therefore we can rewrite (2.3), for $t\approx 0$, as
(2.4) $M(x{\bf 1}+ts)$ $=M(x{\bf 1})+t\sum_{i=1}^{p}\frac{s_{i}}{p}+\frac{t^{2}}{2}\Big{(}\alpha(x)\sum_{i=1}^{p}s_{i}^{2}+\beta(x)\sum_{i\neq j}s_{i}s_{j}\Big{)}+o(t^{2}).$
In a particular case $s=\bf 1$, this identity yields
$x+t=M(x{\bf 1}+t{\bf 1})$ $=x+t+\tfrac{t^{2}}{2}\big{(}p\alpha(x)+(p^{2}-p)\beta(x)\big{)}+o(t^{2}).$
If we subtract $(x+t)$ side-by-side, divide by $t^{2}$ and take the limit as $t\to 0$, we arrive at $p\alpha(x)+(p^{2}-p)\beta(x)=0$. After easy simplification, we obtain $\alpha(x)=(1-p)\beta(x)$.
Furthermore, using the standard identity for symmetric polynomials of degree two, we get
$\sum_{i\neq j}s_{i}s_{j}$ $=\Big{(}\sum_{i=1}^{p}s_{i}\Big{)}^{2}-\sum_{i=1}^{p}s_{i}^{2}=p^{2}({\bf E}s)^{2}-p{\bf E}s^{2}.$
Thus, by (2.4), we obtain
(2.5) $M(x{\bf 1}+ts)$ $=x+t{\bf E}s+\frac{t^{2}}{2}\left(\alpha(x)p{\bf E}s^{2}+\beta(x)\big{(}p^{2}({\bf E}s)^{2}-p{\bf E}s^{2}\big{)}\right)+o(t^{2}).$
Now, let us simplify
$\alpha(x)p{\bf E}s^{2}+\beta(x)\big{(}p^{2}({\bf E}s)^{2}-p{\bf E}s^{2}\big{)}$ $=p(1-p)\beta(x){\bf E}s^{2}+\beta(x)\big{(}p^{2}({\bf E}s)^{2}-p{\bf E}s^{2}\big{)}$
$=\big{(}(p(1-p)-p){\bf E}s^{2}+p^{2}({\bf E}s)^{2}\big{)}\beta(x)$
$=-p^{2}\beta(x)\big{(}{\bf E}s^{2}-({\bf E}s)^{2}\big{)}.$
Then equality above jointly with (2.5) implies (2.1) with the substitution $\xi_{M}(x):=-p^{2}\beta(x)$. Applying definition of $\beta$, we obtain $\xi_{M}(x)=-p^{2}\partial_{1}\partial_{2}M(x{\bf 1})$ which shows the first equality in (2.2). Similarly
$\xi_{M}(x)=-p^{2}\beta(x)=\frac{-p^{2}}{1-p}\alpha(x)=\frac{p^{2}}{p-1}\partial_{1}^{2}M(x{\bf 1}),$
which is the second form of the function $\xi_{M}$ appearing in (2.2). ∎
Based on the above lemma, for every symmetric mean $M\colon I^{p}\to I$ which is $\mathcal{C}^{2}$ in the neighborhood of $\Delta_{p}(I)$ the function $\xi_{M}\colon I\to I$ defined by (2.2) is called the residuum of $M$.
In the following technical lemma, we show a bit easier way to calculate the value of residuum.
Lemma 2.2.
Let $M\colon I^{p}\to I$ be a symmetric mean of class $\mathcal{C}^{2}$ in the neighborhood of the diagonal $\Delta_{p}(I)$. Then
$\xi_{M}(x)=\lim_{\varepsilon\to 0}\frac{2p^{2}}{(p-1)\varepsilon^{2}}\big{(}M(x+\varepsilon,\underbrace{x,\dots,x}_{(p-1)\text{entries}})-x-\tfrac{\varepsilon}{p}\big{)}\qquad(x\in I).$
Proof.
Define $s:=(1,0,0,\dots)$. Then $\operatorname{Var}(s)=\frac{p-1}{p^{2}}$. Thus, by Lemma 2.1 we get
$\xi_{M}(x)=\lim_{\varepsilon\to 0}\frac{2\big{(}M(x{\bf 1}+\varepsilon s)-x-\varepsilon{\bf E}s\big{)}}{\varepsilon^{2}\operatorname{Var}(s)}=\lim_{\varepsilon\to 0}\frac{2p^{2}}{p-1}\frac{M(x{\bf 1}+\varepsilon s)-x-\frac{\varepsilon}{p}}{\varepsilon^{2}},$
which completes the proof. ∎
2.1. Residual means
We say that a mean $M\colon I^{p}\to I$ is residual if for every compact subinterval $J\Subset I$ (which means that the closure of $J$ is contained in the interior of $I$) there exist $K,\varepsilon\in(0,+\infty)$ and $\alpha\in(2,+\infty)$ such that
$\left|M(x{\bf 1}+s)-x-{\bf E}s-\tfrac{1}{2}\xi_{M}(x)\operatorname{Var}(s)\right|<K\left\|s\right\|^{\alpha}$
for all $x\in J$ and $s\in(-\varepsilon,\varepsilon)^{p}$, where $\xi_{M}$ is given by (2.2). Obviously, this property does not depend on the choice of the norm. For the technical purpose assume that this norm is fixed (but arbitrary) and let $S^{p-1}\subset\mathbb{R}^{p}$ be the sphere.
Let $\mathcal{R}_{p}(I)$ be a family of all residual $p$-variable means on $I$. It is obvious that every residual mean is twice continuously differentiable at every point of the diagonal $\Delta_{p}(I)$. We show a sort of reverse statement.
Theorem 2.3.
Every symmetric mean $M\colon I^{p}\to I$ which is $\mathcal{C}^{3}$ in some open neighborhood of $\Delta_{p}(I)$ is residual.
Proof.
Fix $M$ as in the statement, a compact interval $J\Subset I$ and choose $\varepsilon>0$ such that $[\inf J-2\varepsilon,\sup J+2\varepsilon]\subset I$.
To prove that $M$ is residual, it is sufficient to show that there exists $K\in(0,+\infty)$ such that
(2.6) $\left|M(x{\bf 1}+ts)-x-t{\bf E}s-\tfrac{t^{2}}{2}\xi_{M}(x)\operatorname{Var}(s)\right|<K\left|t\right|^{3}$
holds for all $x\in J$, $s\in S^{p-1}$ and $t\in(0,\varepsilon)$. Indeed, this property implies that for all $x\in J$ and a nonzero vector $s^{*}$ with $\left\|s^{*}\right\|<\varepsilon$ we have
$\big{|}M(x{\bf 1}+s^{*})$ $-x-{\bf E}s^{*}-\tfrac{1}{2}\xi_{M}(x)\operatorname{Var}(s^{*})\big{|}$
$=\left|M(x{\bf 1}+\left\|s^{*}\right\|s)-x-\left\|s^{*}\right\|{\bf E}s-\tfrac{\left\|s^{*}\right\|^{2}}{2}\xi_{M}(x)\operatorname{Var}(s)\right|<K\left\|s^{*}\right\|^{3},$
where $s:=\frac{s^{*}}{\left\|s^{*}\right\|}\in S^{p-1}$ is the normalized vector $s^{*}$.
In view of Taylor’s theorem, there exist continuous functions $A,B\colon J\times S^{p-1}\to\mathbb{R}$ and $C\colon J\times S^{p-1}\times[-\varepsilon,\varepsilon]\to\mathbb{R}$ such that
$M(x{\bf 1}+ts)=M(x{\bf 1})+A(x,s)t+B(x,s)t^{2}+C(x,s,t)t^{3}$
for all $x\in J$, $s\in S^{p-1}$, and $t\in[-\varepsilon,\varepsilon]$. Moreover
$A(x,s)$ $:=\phantom{1}\partial^{\phantom{2}}_{t}M(x{\bf 1}+ts)\big{|}_{t=0};$
$B(x,s)$ $:=\tfrac{1}{2}\partial^{2}_{t}M(x{\bf 1}+ts)\big{|}_{t=0};$
$C(x,s,r)$ $\phantom{:}=\tfrac{1}{6}\partial^{3}_{t}M(x{\bf 1}+ts)\big{|}_{t=t^{*}}\text{ for some }t^{*}\in[-\varepsilon,\varepsilon]\qquad(r\in[-\varepsilon,\varepsilon]).$
Since $M$ is $\mathcal{C}^{3}$, we know that the mapping
$C_{0}\colon J\times S^{p-1}\times[-\varepsilon,\varepsilon]\ni(x,s,t_{0})\to\tfrac{1}{6}\partial^{3}_{t}M(x{\bf 1}+ts)|_{t=t_{0}}$
is continuous and, since it is defined on a compact set, bounded. Set $K:=\sup\left|C_{0}\right|\in[0,+\infty)$. Then we also have $\sup\left|C\right|\leq\sup\left|C_{0}\right|=K$.
In the next step, let us define the reminder $R\colon J\times S^{p-1}\times[-\varepsilon,\varepsilon]\to\mathbb{R}$ by
$R(x,s,t):=M(x{\bf 1}+ts)-M(x{\bf 1})-A(x,s)t-B(x,s)t^{2}=C(x,s,t)t^{3}.$
Then, obviously, $\left|R(x,s,t)\right|\leq K\left|t\right|^{3}$ (in the domain of $R$). In order to complete the proof of (2.6), it is sufficient to show that
$R(x,s,t)$ $=M(x{\bf 1}+ts)-x-t{\bf E}s-\tfrac{t^{2}}{2}\xi_{M}(x)\operatorname{Var}(s),$
which is the straightforward implication of Lemma 2.1 and Taylor’s theorem. As we already mentioned, inequality (2.6) implies that $M$ is residual. ∎
2.2. Means on any number of arguments
Until now all means considered in this note were defined only for a given (fixed) number of arguments. It is, however, quite often that the input to the mean can be a vector of an arbitrary (finite) length. Then a mean on $I$ is a function $M\colon\bigcup_{p=1}^{\infty}I^{p}\to I$ such that each $M\!\!\upharpoonright_{p}:=M|_{I^{p}}$ is a $p$-variable mean on $I$. Then we use the standard convention that such a mean has a certain property (for example, continuity, symmetry, etc.) precisely if all its restrictions $M\!\!\upharpoonright_{p}$ admit this property. In the same spirit, we say that a mean $M\colon\bigcup_{p=1}^{\infty}I^{p}\to I$ is residual if, for all $p\in\mathbb{N}$, its $p$-variable restrictions $M\!\!\upharpoonright_{p}$ are residual.
We also need one property which is of different nature. Namely, mean $M$ is called repetition invariant if, for all $p,q\in\mathbb{N}$ and $(x_{1},\dots,x_{p})\in I^{p}$, the following identity is satisfied
$M(\underbrace{x_{1},\dots,x_{1}}_{q\text{-times}},\dots,\underbrace{x_{p},\dots,x_{p}}_{q\text{-times}})=M(x_{1},\dots,x_{p}).$
This definition was first introduced by Páles-Pasteczka [37]. It turns out that this property holds for a number of classical families of means – for example, a broad class of quasideviation means (see section 4.1 below).
In the next lemma, we show that this property also plays an important role in our new concept of residual means.
Lemma 2.4.
Let $M\colon\bigcup_{p=1}^{\infty}I^{p}\to I$ be a symmetric, repetition invariant, residual mean. Then the value of $\xi_{M\,\!\!\upharpoonright_{p}}$ does not depend on $p$.
Proof.
Let $p,q\in\mathbb{N}$, $p,q\geq 2$. Instead of proving the equality $\xi_{M\,\!\!\upharpoonright_{p}}=\xi_{M\,\!\!\upharpoonright_{q}}$, we show that $\xi_{M\,\!\!\upharpoonright_{pq}}=\xi_{M\,\!\!\upharpoonright_{p}}$, since these two properties are equivalent (provided they hold for all $p,q\in\mathbb{N}\setminus\{1\}$). To keep the proof more compact, fix $x\in I$ and define
$\hat{x}:=(\underbrace{x,\dots,x}_{p\text{ times}}),\quad\hat{s}:=(1,\underbrace{0,\dots,0}_{(p-1)\text{ times}}),\quad\bar{x}:=(\underbrace{x,\dots,x}_{pq\text{ times}}),\quad\bar{s}:=(\underbrace{1,\dots,1}_{q\text{ times}},\underbrace{0,\dots,0}_{(pq-q)\text{ times}}).$
Then $\bar{s}$ and $\hat{s}$ possess the same probability distribution (provided choosing entries are uniformly distributed), and whence ${\bf E}\hat{s}={\bf E}\bar{s}$, and $\operatorname{Var}(\hat{s})=\operatorname{Var}(\bar{s})$. Furthermore, since $M$ is repetition invariant, we have $M(\bar{x}+t\bar{s})=M(\hat{x}+t\hat{s})$ for all admissible $t$. Thus, Lemma 2.2 implies (here and below $t\approx 0$)
00 0 $=M(\bar{x}+t\bar{s})-M(\hat{x}+t\hat{s})$
$=x+t{\bf E}\bar{s}+\xi_{M\,\!\!\upharpoonright_{pq}}(x)\operatorname{Var}(\bar{s})\tfrac{t^{2}}{2}-x-t{\bf E}\hat{s}-\xi_{M\,\!\!\upharpoonright_{p}}(x)\operatorname{Var}(\hat{s})\tfrac{t^{2}}{2}+o(t^{2})$
$=\tfrac{t^{2}}{2}\operatorname{Var}(\hat{s})(\xi_{M\,\!\!\upharpoonright_{pq}}(x)-\xi_{M\,\!\!\upharpoonright_{p}}(x))+o(t^{2}).$
Since $\hat{s}$ is a nonconstant vector we know that $\operatorname{Var}(\hat{s})\neq 0$. Thus, we can divide by $\frac{t^{2}}{2}\operatorname{Var}(\hat{s})$ and take the limit $t\to 0$, to obtain $0=\xi_{M\,\!\!\upharpoonright_{pq}}(x)-\xi_{M\,\!\!\upharpoonright_{p}}(x)$. Since $x$ was chosen to be an arbitrary element in $I$, we immediately get $\xi_{M\,\!\!\upharpoonright_{pq}}=\xi_{M\,\!\!\upharpoonright_{p}}$. This equality completes the proof, as it was announced at its beginning. ∎
3. An application to invariant means
A mapping $\mathbf{M}\colon I^{p}\rightarrow I^{p}$ is referred to as mean-type if $\mathbf{M}=(M_{1},\dots,M_{p})$ for some means $M_{1},\dots,M_{p}\colon I^{p}\rightarrow I$. A mean $K\colon I^{p}\rightarrow I$ is invariant with respect to the mean-type mapping $\mathbf{M}$ (briefly $\mathbf{M}$-invariant), if it solves the functional equation $K\circ\mathbf{M}=K$.
Invariance property is a vital aspect of the theory of means. There are two classical studies, Lagrange [22] and Gauss [14], which could be considered as the beginning of this field. It has been extensively studied by many authors since then. For example J. M. Borwein and P. B. Borwein [7] extended some earlier ideas [13, 23, 39] and generalized the original iteration to a vector of continuous, strict means of an arbitrary length. For several recent results about the Gaussian product of means, see the papers by Baják–Páles [3, 4, 5, 6], by Daróczy–Páles [9, 10, 11], by Głazowska [16, 17], by Jarczyk–Jarczyk [19], by Matkowski [24, 25, 26, 27], by Matkowski-Pasteczka [28, 29], by Matkowski–Páles [30], and by Pasteczka [33, 35, 34].
There are very close relations between the invariant means and the sequence of iterations of $\mathbf{M}$. For example, due to [29], the $\mathbf{M}$-invariant mean is uniquely determined if, and only if, for all $x\in I^{p}$ the sequence of iterates $(\mathbf{M}^{n}(x))_{n=1}^{\infty}$ is convergent to some point of $\Delta_{p}(I)$. On the other hand, in Borwein-Borwein [7, Theorem 8.8], it was proved that if means are twice continuously differentiable, then convergence in this iteration is quadratic. In this spirit, the author proved [33] that in the case when the mean-type mapping consists of quasiarithmetic means satisfying some additional smoothness assumptions, we can efficiently calculate the limit of the sequence of ratios $\Big{(}\frac{\operatorname{Var}\mathbf{M}^{n+1}(x)}{\operatorname{Var}\mathbf{M}^{n}(x)^{2}}\Big{)}_{n=1}^{\infty}$ in the nondegenerated case; see Theorem A above. The limit value turned out to be the variance of Arrow-Pratt indexes of the corresponding means calculated at the point that is equal to the invariant mean. It will be mostly generalized in our next theorem. For the sake of completeness, let us mention that there were slightly different smoothness assumptions in the previous result.
Similarly to Theorem A, we show that if the mean-type mapping $\mathbf{M}$ contains residual means only, then the sequence $\operatorname{Var}\mathbf{M}^{n}(x)$ of iterations of the vector converges to zero quadratically for >>generic<< vectors $x$ and we calculate the precise speed of this convergence.
Theorem 3.1.
Let $p\in\mathbb{N}$, and $\mathbf{M}:=(M_{1},\dots,M_{p})\colon I^{p}\to I^{p}$ be a mean-type mapping consisting of residual means. If there exists a unique $\mathbf{M}$-invariant mean $K\colon I^{p}\to I$ then for each $x\in I^{p}$ either there exists $n_{0}\in\mathbb{N}$ such that $\mathbf{M}^{n_{0}}(x)$ is constant or
$\lim_{n\to\infty}\frac{\operatorname{Var}\mathbf{M}^{n+1}(x)}{(\operatorname{Var}\mathbf{M}^{n}(x))^{2}}=\tfrac{1}{4}\operatorname{Var}\big{(}\xi_{M_{1}}\big{(}K(x)\big{)},\dots,\xi_{M_{p}}\big{(}K(x)\big{)}\big{)}.$
Proof.
Let $x\in I^{p}$. Define the sequence $(y^{(n)})_{n=0}^{\infty}$ of the elements in $I^{p}$ by $y^{(0)}:=x$, $y^{(n)}:=\mathbf{M}^{n}(x)$. Moreover, let $\mu:=K(x)$, and $m_{n}$ be the arithmetic mean of elements in $y^{(n)}$ ($n\in\mathbb{N}$). Then, since $K$ is $\mathbf{M}$-invariant, we get $\mu=K(y^{(n)})$ for all $n\in\mathbb{N}$.
In order to show this statement, we can claim that all $y^{(n)}$-s are nonconstant vectors. Then
$\operatorname{Var}(y^{(n+1)})=\operatorname{Var}\Big{(}\big{(}M_{i}(y^{(n)})\big{)}_{i=1}^{p}\Big{)}=\operatorname{Var}\Big{(}\big{(}M_{i}(y^{(n)})-m_{n}\big{)}_{i=1}^{p}\Big{)}.$
But, since all $M_{i}$-s are residual, there exist $C>0$ and $\alpha>2$ such that
$y^{(n+1)}_{i}=M_{i}(y^{(n)})=m_{n}+\frac{\xi_{M_{i}}(\mu)}{2}\operatorname{Var}(y^{(n)})\pm e_{n}\text{ for all }n\in\mathbb{N}\text{ and }i\in\{1,\dots,p\},$
where $e_{n}:=C\sum_{k=1}^{p}\left|y^{(n)}_{k}-\mu\right|^{\alpha}$. Consequently
$m_{n+1}=\frac{1}{p}\sum_{i=1}^{p}y^{(n+1)}_{i}=m_{n}+\frac{1}{2p}\sum_{i=1}^{p}\xi_{M_{i}}(\mu)\operatorname{Var}(y^{(n)})\pm e_{n}.$
Thus, for all $i\in\{1,\dots,p\}$, we have
$\Big{|}y^{(n+1)}_{i}$ $-m_{n+1}\Big{|}$
$=\left|m_{n}+\frac{1}{2}\xi_{M_{i}}(\mu)\operatorname{Var}(y^{(n)})\pm e_{n}-\Big{(}m_{n}+\frac{1}{2p}\sum_{j=1}^{p}\xi_{M_{j}}(\mu)\operatorname{Var}(y^{(n)})\pm e_{n}\Big{)}\right|$
$=\left|\frac{1}{2}\xi_{M_{i}}(\mu)\operatorname{Var}(y^{(n)})-\frac{1}{2p}\sum_{j=1}^{p}\xi_{M_{j}}(\mu)\operatorname{Var}(y^{(n)})\pm 2e_{n}\right|$
$=\left|\frac{1}{2}\xi_{M_{i}}(\mu)-\frac{1}{2p}\sum_{j=1}^{p}\xi_{M_{j}}(\mu)\pm\frac{2e_{n}}{\operatorname{Var}(y^{(n)})}\right|\operatorname{Var}(y^{(n)}).$
Therefore
$\operatorname{Var}(y^{(n+1)})$ $=\frac{1}{p}\sum_{i=1}^{p}(y^{(n+1)}_{i}-m_{n+1})^{2}$
$=\frac{1}{p}\sum_{i=1}^{p}\Bigg{(}\frac{1}{2}\xi_{M_{i}}(\mu)-\frac{1}{2p}\sum_{j=1}^{p}\xi_{M_{j}}(\mu)\pm\frac{2e_{n}}{\operatorname{Var}(y^{(n)})}\Bigg{)}^{2}\operatorname{Var}(y^{(n)})^{2}.$
Whence
(3.1) $\frac{\operatorname{Var}(y^{(n+1)})}{\operatorname{Var}(y^{(n)})^{2}}$ $=\frac{1}{4p}\sum_{i=1}^{p}\Bigg{(}\xi_{M_{i}}(\mu)-\frac{1}{p}\sum_{j=1}^{p}\xi_{M_{j}}(\mu)\pm\frac{2e_{n}}{\operatorname{Var}(y^{(n)})}\Bigg{)}^{2}.$
But
$\frac{2e_{n}}{\operatorname{Var}(y^{(n)})}$ $=\frac{2C\sum_{k=1}^{p}\left|y^{(n)}_{k}-m_{n}\right|^{\alpha}}{\sum_{k=1}^{p}\left|y^{(n)}_{k}-m_{n}\right|^{2}}\leq 2C\sup_{k\in\{1,\dots,p\}}\left|y^{(n)}_{k}-m_{n}\right|^{\alpha-2}$
$\leq 2C(\max y^{(n)}-\min y^{(n)})^{\alpha-2}$
Since $\alpha>2$ and $\lim_{n\to\infty}\max y^{(n)}-\min y^{(n)}=0$, we get
$\lim_{n\to\infty}\frac{2e_{n}}{\operatorname{Var}(y^{(n)})}=0.$
Then (3.1) yields
$\lim_{n\to\infty}\frac{\operatorname{Var}(y^{(n+1)})}{\operatorname{Var}(y^{(n)})^{2}}$ $=\frac{1}{4p}\sum_{i=1}^{p}\Big{(}\xi_{M_{i}}(\mu)-\frac{1}{p}\sum_{j=1}^{p}\xi_{M_{j}}(\mu)\Big{)}^{2}$
$=\tfrac{1}{4}\operatorname{Var}\big{(}\xi_{M_{1}}(\mu),\dots,\xi_{M_{p}}(\mu)\big{)}$
$=\tfrac{1}{4}\operatorname{Var}\big{(}\xi_{M_{1}}\big{(}K(x)\big{)},\dots,\xi_{M_{p}}\big{(}K(x)\big{)}\big{)},$
which was to be proved. ∎
4. Examples
The aim of this section is to establish residua for possibly broad classes of means. This allows one to apply Theorem 3.1 for all mean-type mapping built from means in these classes. We intentionally avoid direct applications of Theorem 3.1, since it is a purely technical operation.
4.1. Quasideviation means
The notions of deviations and quasideviations were introduced by Daróczy in [8] and by Páles [36], respectively. In what follows, we recall Definition 2.1 and Theorem 2.1 from [36]. A bivariate function $E:I^{2}\to\mathbb{R}$ will be called a quasideviation if $E$ possesses the following three properties:
  1. (D1)
    For all $x,u\in I$, the equality $\operatorname{sign}E(x,u)=\operatorname{sign}(x-u)$ holds;
  2. (D2)
    For all $x\in I$, the mapping $I\ni u\mapsto E(x,u)$ is continuous;
  3. (D3)
    For all $x,y\in I$ with $x<y$, the mapping
    $(x,y)\ni u\mapsto\frac{E(x,u)}{E(y,u)}$
    is strictly decreasing.
In order to introduce quasideviation means, the following statement is instrumental (cf. [36, Theorem 2.1]).
Proposition 4.1.
Let $E:I^{2}\to\mathbb{R}$ be a quasideviation. Then, for all $n\in\mathbb{N}$ and $x_{1},\dots,x_{n}\in I$, there exists a unique element $u\in I$ such that
(4.1) $E(x_{1},u)+\dots+E(x_{n},u)=0.$
Furthermore, $\min(x_{1},\dots,x_{n})<u<\max(x_{1},\dots,x_{n})$ unless $x_{1}=\dots=x_{n}$.
For $n\in\mathbb{N}$ and $x_{1},\dots,x_{n}\in I$, the solution $u$ of equation (4.1) is called the $E$-quasideviation mean of $x_{1},\dots,x_{n}$ and will be denoted by $\mathscr{D}_{E}(x_{1},\dots,x_{n})$. It is easy to check that quasideviation means are repetition invariant.
We say that a quasideviation $E\colon I\times I\to\mathbb{R}$ is normalizable (cf. [8]) if, for all $x\in I$, the function $u\mapsto E(x,u)$ is differentiable at $x$ and the mapping $x\mapsto\partial_{2}E(x,x)$ is strictly negative and continuous on $I$.
Theorem 4.2.
Let $E\in\mathcal{C}^{2}(I^{2})$ be a normalizable quasideviation. Then $\mathscr{D}_{E}$ is twice continuously differentiable. Moreover
$\xi_{\mathscr{D}_{E}}(x)=\frac{\partial_{1}^{2}E}{\partial_{1}E}(x,x),\qquad x\in I.$
Proof.
By the definition of quasideviation mean, in view of the implicit function theorem, we know that if $E$ is a normalizable quasideviation of class $\mathcal{C}^{2}$ then $\mathscr{D}_{E}\in\mathcal{C}^{2}$.
Fix $p\in\mathbb{N}$, $x\in I$ and take any $\varepsilon_{0}>0$ such that $(x-\varepsilon_{0},x+\varepsilon_{0})\subset I$. For all $\varepsilon\in(-\varepsilon_{0},\varepsilon_{0})\setminus\{0\}$ we define
$\delta=\delta(\varepsilon):=\mathscr{D}_{E}(x+\varepsilon,\underbrace{x,\dots,x}_{(p-1)\text{entries}})-x.$
Then $\delta$ is the unique solution of the equation
00 0 $=E(x+\varepsilon,x+\delta)+(p-1)E(x,x+\delta)$
$=\partial_{1}E(x,x)\varepsilon+\partial_{2}E(x,x)\delta+\tfrac{1}{2}\partial_{1}^{2}E(x,x)\varepsilon^{2}+\tfrac{1}{2}\partial_{2}^{2}E(x,x)\delta^{2}+\partial_{1}\partial_{2}E(x,x)\varepsilon\delta$
$\quad+(p-1)\big{(}\partial_{2}E(x,x)\delta+\tfrac{1}{2}\partial_{2}^{2}E(x,x)\delta^{2}\big{)}.$
But property (D1) implies $\partial_{1}E(x,x)=-\partial_{2}E(x,x)$, and whence
(4.2) $0=\big{(}\varepsilon-p\delta\big{)}\partial_{1}E(x,x)+\tfrac{1}{2}\big{(}\partial_{1}^{2}E(x,x)\varepsilon^{2}+2\partial_{1}\partial_{2}E(x,x)\varepsilon\delta+p\partial_{2}^{2}E(x,x)\delta^{2}\big{)}$
Now define $c\colon(-\varepsilon_{0},\varepsilon_{0})\setminus\{0\}\to\mathbb{R}$ by $c(\varepsilon):=\frac{1}{p\varepsilon^{2}}(p\delta(\varepsilon)-\varepsilon)$. Then we have
(4.3) $\delta(\varepsilon)=\tfrac{\varepsilon}{p}+\varepsilon^{2}c(\varepsilon).$
Whence, by Lemma 2.2, we have
$\xi_{\mathscr{D}_{E}}(x)$ $=\lim_{\varepsilon\to 0}\frac{2p^{2}}{(p-1)\varepsilon^{2}}(\mathscr{D}_{E}(x+\varepsilon,\underbrace{x,\dots,x}_{(p-1)\text{entries}})-x-\tfrac{\varepsilon}{p})$
$=\frac{2p^{2}}{p-1}\lim_{\varepsilon\to 0}\dfrac{1}{\varepsilon^{2}}(\delta(\varepsilon)-\tfrac{\varepsilon}{p})=\frac{2p^{2}}{p-1}\lim_{\varepsilon\to 0}c(\varepsilon).$
Thus $c$ has a finite limit at 00, say $c_{0}\in\mathbb{R}$, and $\xi_{\mathscr{D}_{E}}(x)=\frac{2p^{2}}{p-1}c_{0}$. We now substitute equality (4.3) to (4.2) and obtain
$p\varepsilon^{2}c(\varepsilon)\partial_{1}E(x,x)$ $=\frac{\varepsilon^{2}}{2}\Big{(}\partial_{1}^{2}E(x,x)+2\partial_{1}\partial_{2}E(x,x)\big{(}\tfrac{1}{p}+\varepsilon c(\varepsilon)\big{)}+p\partial_{2}^{2}E(x,x)\big{(}\tfrac{1}{p}+\varepsilon c(\varepsilon)\big{)}^{2}\Big{)}$
$c(\varepsilon)$ $=\frac{1}{2p}\frac{\partial_{1}^{2}E(x,x)}{\partial_{1}E(x,x)}+\frac{\partial_{1}\partial_{2}E(x,x)}{{\partial_{1}E(x,x)}}\frac{\tfrac{1}{p}+\varepsilon c(\varepsilon)}{p}+\frac{\partial_{2}^{2}E(x,x)}{\partial_{1}E(x,x)}\frac{\big{(}\tfrac{1}{p}+\varepsilon c(\varepsilon)\big{)}^{2}}{2}.$
Upon passing the limit $\varepsilon\to 0$ we have
$c_{0}$ $=\frac{1}{2p}\frac{\partial_{1}^{2}E(x,x)}{\partial_{1}E(x,x)}+\frac{\partial_{1}\partial_{2}E(x,x)}{{\partial_{1}E(x,x)}}\frac{1}{p^{2}}+\frac{\partial_{2}^{2}E(x,x)}{\partial_{1}E(x,x)}\frac{1}{2p^{2}}$
$=\frac{1}{2p^{2}}\bigg{(}\frac{p\partial_{1}^{2}E+2\partial_{1}\partial_{2}E+\partial_{2}^{2}E}{\partial_{1}E}\bigg{)}(x,x).$
Now note that for all $\varepsilon\in(-\varepsilon_{0},\varepsilon_{0})$ we have
$0=E(x+\varepsilon,x+\varepsilon)$ $=E(x,x)+(\partial_{1}E+\partial_{2}E)(x,x)\varepsilon+(\partial_{1}^{2}E+2\partial_{1}\partial_{2}E+\partial_{2}^{2}E)(x,x)\varepsilon^{2}.$
Thus we get
$(\partial_{1}E+\partial_{2}E)(x,x)=0\quad\text{and}\quad(\partial_{1}^{2}E+2\partial_{1}\partial_{2}E+\partial_{2}^{2}E)(x,x)=0\qquad(x\in I).$
Finally, we obtain
$\xi_{\mathscr{D}_{E}}(x)$ $=\frac{2p^{2}}{(p-1)}c_{0}=\bigg{(}\frac{p\partial_{1}^{2}E+2\partial_{1}\partial_{2}E+\partial_{2}^{2}E}{(p-1)\partial_{1}E}\bigg{)}(x,x)$
$=\bigg{(}\frac{p\partial_{1}^{2}E-\partial_{1}^{2}E}{(p-1)\partial_{1}E}\bigg{)}(x,x)=\frac{\partial_{1}^{2}E}{\partial_{1}E}(x,x),$
which completes the proof. ∎
4.2. Bajraktarević means
Now we study an important subclass of quasideviation means defined by Bajraktarević in the papers [1, 2]. Let $f,g\colon I\to\mathbb{R}$ be two functions such that $g>0$ and the ratio $x\mapsto\frac{f(x)}{g(x)}$ is continuous and strictly increasing. Then
(4.4) $E\colon I^{n}\ni(x,u)\mapsto g(u)f(x)-f(u)g(x)$
is quasideviation, and a simple computation yields that
$\mathscr{B}_{f,g}(x):=\mathscr{D}_{E}(x)=\Big{(}\frac{f}{g}\Big{)}^{-1}\left(\frac{f(x_{1})+\cdots+f(x_{p})}{g(x_{1})+\cdots+g(x_{p})}\right)\qquad(x\in I^{p}).$
Before we proceed to calculate the value of residuum let us recall the theorem which characterizes the equality of Bajraktarević means. Namely, for the functions $f,g\in\mathcal{C}^{2}(I)$, Páles-Zakaria [38] defined two auxiliary operators $\Phi_{f,g},\Psi_{f,g}\colon I\to\mathbb{R}$ by
$\Phi_{f,g}(x):=\bigg{(}\frac{gf^{\prime\prime}-fg^{\prime\prime}}{gf^{\prime}-fg^{\prime}}\bigg{)}(x),\qquad\Psi_{f,g}(x):=\bigg{(}\frac{g^{\prime}f^{\prime\prime}-f^{\prime}g^{\prime\prime}}{gf^{\prime}-fg^{\prime}}\bigg{)}(x);$
and proved that for $f,g,h,k\colon I\to\mathbb{R}$ (satisfying certain smoothness assumptions) the equality $\mathscr{B}_{f,g}=\mathscr{B}_{h,k}$ holds if and only if $\Phi_{f,g}=\Phi_{h,k}$ and $\Psi_{f,g}=\Psi_{h,k}$ (cf. [38, Theorem 2.1]).
We show that the function $\Phi_{f,g}$ is the residuum of $\mathscr{B}_{f,g}$. We can combine this fact with the above equality conditions to construct Bajraktarević means which have the same residuum, although they are not equal to each other. Remarkable, if the mean-type mapping consists of means with the same residuum then, in view of Theorem 2.3, the convergence of (variances of) its iterates is superqudratic.
Proposition 4.3.
Let $f,g\in\mathcal{C}^{2}(I)$ such that $g>0$, the ratio $x\mapsto\frac{f(x)}{g(x)}$ is strictly increasing, and the mapping $gf^{\prime}-fg^{\prime}$ is nowhere vanishing. Then $\xi_{\mathscr{B}_{f,g}}=\Phi_{f,g}$.
Proof.
We know that $\mathscr{B}_{f,g}=\mathscr{D}_{E}$, where $E\colon I\times I\to\mathbb{R}$ is of the form (4.4). Moreover
$\partial_{1}E(x,u)=g(u)f^{\prime}(x)-f(u)g^{\prime}(x)\text{ and }\partial_{1}^{2}E(x,u)=g(u)f^{\prime\prime}(x)-f(u)g^{\prime\prime}(x).$
Therefore, by Theorem 4.2, for all $x\in I$, we have
$\xi_{\mathscr{B}_{f,g}}(x)=\xi_{\mathscr{D}_{E}}(x)=\frac{\partial_{1}^{2}E}{\partial_{1}E}(x,x)=\frac{g(x)f^{\prime\prime}(x)-f(x)g^{\prime\prime}(x)}{g(x)f^{\prime}(x)-f(x)g^{\prime}(x)}=\Phi_{f,g}(x),$
which completes the proof. ∎
4.3. Gini means
Now we study a generalization of Hölder means introduced by Gini [15]. This is simultaneously a subclass of Bajraktarević means.
More precisely, for $\alpha,\beta\in\mathbb{R}$, define the Gini mean $\mathscr{G}_{\alpha,\beta}\colon\bigcup_{n=1}^{\infty}\mathbb{R}_{+}^{n}\to\mathbb{R}_{+}$ by
$\mathscr{G}_{\alpha,\beta}(x):=\begin{cases}\left(\dfrac{x_{1}^{\alpha}+\cdots+x_{n}^{\alpha}}{x_{1}^{\beta}+\cdots+x_{n}^{\beta}}\right)^{\frac{1}{\alpha-\beta}}&\quad\text{ if }\alpha\neq\beta,\\[11.38109pt] \exp\left(\dfrac{x_{1}^{\alpha}\log x_{1}+\cdots+x_{n}^{\alpha}\log x_{n}}{x_{1}^{\alpha}+\cdots+x_{n}^{\alpha}}\right)&\quad\text{ if }\alpha=\beta.\end{cases}$
Remarkably, if $\alpha,\beta\in\mathbb{R}$,
$f(x):=\begin{cases}x^{\alpha}&\text{ if }\alpha\neq\beta,\\ x^{\alpha}\log(x)&\text{ if }\alpha=\beta,\end{cases}\qquad\text{ and }\qquad g(x):=x^{\beta},$
then $\mathscr{B}_{f,g}=\mathscr{G}_{\alpha,\beta}$. We can utilize this fact to establish the residuum of all means belonging to this family.
Proposition 4.4.
For all $\alpha,\beta\in\mathbb{R}$ we have $\xi_{\mathscr{G}_{\alpha,\beta}}(x)=\frac{\alpha+\beta-1}{x}$.
Proof.
In view of Proposition 4.3, for $\alpha\neq\beta$ and $x\in\mathbb{R}_{+}$ we have
$\xi_{\mathscr{G}_{\alpha,\beta}}(x)$ $=\frac{x^{\beta}(x^{\alpha})^{\prime\prime}-x^{\alpha}(x^{\beta})^{\prime\prime}}{(x^{\beta}(x^{\alpha})^{\prime}-x^{\alpha}(x^{\beta})^{\prime}}=\frac{(\alpha(\alpha-1)-\beta(\beta-1))x^{\alpha+\beta-2}}{(\alpha-\beta)x^{\alpha+\beta-1}}$
$=\frac{(\alpha-\beta)(\alpha+\beta-1)}{(\alpha-\beta)x}=\frac{\alpha+\beta-1}{x}.$
Similarly, in the case $\alpha=\beta$, we obtain $\xi_{\mathscr{G}_{\alpha,\alpha}}(x)=\frac{2\alpha-1}{x}=\frac{\alpha+\beta-1}{x}$. ∎
Remark.
Observe that, in view of the above proposition, we can construct the family of (pairwise) distinct Gini means with the same residuum.
4.4. Quasiarithmetic means
Another important generalization of Hölder’s means is the notion of quasiarithmetic means. This family was introduced in the series of several simultaneous papers [20, 12, 21, 32] in 1920-s/30-s as a generalization of power means. Given a continuous strictly monotone function $f\colon I\to\mathbb{R}$, the quasiarithmetic mean $\mathscr{A}_{f}:\bigcup_{n=1}^{\infty}I^{n}\to I$ is defined by
(4.5) $\mathscr{A}_{f}(x):=f^{-1}\left(\frac{f(x_{1})+\cdots+f(x_{n})}{n}\right).$
If $\alpha\in\mathbb{R}\setminus\{0\}$ and $f(x):=x^{\alpha}$ for $x\in\mathbb{R}_{+}$, then $\mathscr{A}_{f}=\mathscr{H}_{\alpha}$. If $f(x):=\log x$ for $x\in\mathbb{R}_{+}$, then $\mathscr{A}_{f}=\mathscr{H}_{0}$, therefore, Hölder means are quasiarithmetic means.
If $g$ is a constant function, then one can see that $\mathscr{B}_{f,g}=\mathscr{A}_{f}$ and hence quasiarithmetic means are also included in the class of Bajraktarević means. Then, as an immediate result of Proposition 4.3 we can establish the following statement.
Proposition 4.5.
Let $f\in\mathcal{C}^{2}(I)$ be a continuous and strictly increasing function with nowhere vanishing first derivative. Then $\mathscr{A}_{f}$ is twice continuously differentiable, and
$\xi_{\mathscr{A}_{f}}(x)=\frac{f^{\prime\prime}(x)}{f^{\prime}(x)},\qquad x\in I.$
In view of the latter proposition we see that, in the case when all means $M_{i}$ are quasiarithmetic means, Theorem 3.1 simplifies to Theorem A (up to certain issues related to the smoothness assumptions).
4.5. An example
In the last section we show a classical application of our statement in the easy example. Let $\mathbf{M}\colon\mathbb{R}_{+}^{3}\to\mathbb{R}_{+}^{3}$ be given by
$\mathbf{M}(x_{1},x_{2},x_{3}):=\bigg{(}\frac{x_{1}^{2}+x_{2}^{2}+x_{3}^{2}}{x_{1}+x_{2}+x_{3}},\sqrt{x_{1}x_{2}x_{3}},\sqrt{\frac{x_{1}+x_{2}+x_{3}}{x_{1}^{-1}+x_{2}^{-1}+x_{3}^{-1}}}\bigg{)}.$
Then $\mathbf{M}=(\mathscr{G}_{2,1},\mathscr{G}_{0,0},\mathscr{G}_{1,-1})$. Using the classical result concerning the comparability of Gini means, one can show that $\mathbf{M}^{n}(x)$ is nonconstant for every vector $x\in\mathbb{R}_{+}^{3}$ and $n\in\mathbb{N}$. Moreover
$\xi_{\mathscr{G}_{2,1}}(t)=\tfrac{2}{t},\xi_{\mathscr{G}_{0,0}}(t)=-\tfrac{1}{t},\xi_{\mathscr{G}_{1,-1}}(t)=-\tfrac{1}{t}\qquad(t\in\mathbb{R}_{+}).$
Whence $\operatorname{Var}\big{(}\xi_{\mathscr{G}_{2,1}}(t),\xi_{\mathscr{G}_{0,0}}(t),\xi_{\mathscr{G}_{1,-1}}(t)\big{)}=\tfrac{\operatorname{Var}(2,-1,-1)}{t^{2}}=\tfrac{2}{t^{2}}.$ Thus, by Theorem 3.1, we have
$\lim_{n\to\infty}\frac{\operatorname{Var}\mathbf{M}^{n+1}(x)}{(\operatorname{Var}\mathbf{M}^{n}(x))^{2}}=\tfrac{1}{4}\operatorname{Var}\big{(}\xi_{M_{1}}\big{(}K(x)\big{)},\dots,\xi_{M_{p}}\big{(}K(x)\big{)}\big{)}=\frac{1}{2(K(x))^{2}},$
where $K\colon\mathbb{R}_{+}^{3}\to\mathbb{R}_{+}$ is the (unique) $\mathbf{M}$-invariant mean.